a mother invest $13,000.00 for her son's future education she invests a portion of the money in a bank certificate of deposit (cd account) which earns 4% and the remainder in a savings bond that earns 7%. if the total interest after one year is $780.00, how much money was invested in the cd account.
Answer by richwmiller(17219)

x+y=13000,
0.04*x+0.07*y=780
x=13000-y
0.04*(13000-y)+0.07*y=780
520-0.04y+0.07*y=780
0.03*y=260
y=8666.66667 at 7%
x=13000-y
x=4333.33333 at 4%
check
0.04*4333.33333+0.07*8666.66667=780
173.333333+606.666667=780
780.0=780
ok